-8x+3y=-33
5x+y=35
Solve?

Answer by mananth(16946)

-8x+3y =-33.............1
5x+1y=35.............2

multiply (1)by 1
Multiply (2) by -3
-8x+3y =-33
-15x+-3y=-105

Add the two equations

-23x=-138

/-23

x=6

plug value of x in (1)

-8x+3y=-33
-48+3y=-33
3y=-33 +48
3y=15

y=5